OpenStudy (anonymous):

hey can anyone help me with my study gudie please befor 12 i got 2 to do i have got to them becuase somone died in my family so pleaseeee What is the slope of a line parallel to the line with equation 5x – 2y = 11? (Points : 4) 4a.gif 972979_4b.gif 972980_4c.gif 4D.gif Question 5. 5. Which of the following has a slope of 2 and a y-intercept of –4? (Points : 4) y = –2x + 4 y = 2x + 4 y = 2x – 4 y = –2x – 4 Question 6. 6. Which ordered pair is on the graph of the equation 3x + 7y = 11? (Points : 4) (1, –2)

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

\[2x-2y=11\] lets solve for \(y\)

OpenStudy (anonymous):

\[5x-2y=11\\ -2y=-5x+11\\ y=\frac{5}{2}-\frac{11}{2}\]

OpenStudy (anonymous):

can you read off the slope of \[y=\frac{5}{2}-\frac{11}{2}\]?

OpenStudy (anonymous):

oops i meant \[y=\frac{5}{2}x-\frac{11}{2}\]

OpenStudy (anonymous):

5 over 2 neagtive

OpenStudy (anonymous):

no just five over two, i.e. \(\frac{5}{2}\)

OpenStudy (anonymous):

oh okay so the next one would be c

OpenStudy (anonymous):

y = 2x – 4

OpenStudy (anonymous):

yes C

OpenStudy (anonymous):

-4 i mean

OpenStudy (anonymous):

\[y=2x-4\] is the right answer yes

OpenStudy (anonymous):

Which shows solutions to x + 2y = 8 if x and y must be whole numbers. (Points : 4) {(2, 4), (1, 6), (0, 8)} {(8, 0), (6, 1), (4, 2)} {(0, 8), (1, 5), (2, 4)} {(0, 8), (1, 7), (2, 4)

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

you have to plug in all the numbers and see which ones work

OpenStudy (anonymous):

okay i can do on the calc can you help me with 2 more questions i will let you go

OpenStudy (anonymous):

the first coordinate for \(x\) and the second one for \(y\) fortunately you can forget about the last two because if \(x=0, y =8\) you get \[0+2\times 8=8\] which is false it has to be one of the first two

OpenStudy (anonymous):

i think only this one \[ \{(8, 0), (6, 1), (4, 2)\}\] works

OpenStudy (anonymous):

oh okay i just didnt know how to do it thanks Which ordered pair is on the graph of the equation 3x + 7y = 11?

OpenStudy (anonymous):

same question, different line

OpenStudy (anonymous):

oh okay well can yo explain this 2

OpenStudy (anonymous):

Which is the graph of the line with equation: y – 6 = 4(x – 3) + 2

OpenStudy (anonymous):

What is the slope of a line perpendicular to the line with equation y = 6x + 2? (Points : 4) this one i just dont know about the slope wither is negative or positive

OpenStudy (anonymous):

hold on i got the dreaded question marks maybe if i refresh

OpenStudy (anonymous):

this one What is the slope of a line perpendicular to the line with equation y = 6x + 2? we can do in a hearbeat the slope of \(y=6x+2\) is \(6\) the perpendicular line with have a slope that is the "negative reciprocal' i.e. \[-\frac{1}{6}\]

OpenStudy (anonymous):

Which is the graph of the line with equation : y – 6 = 4(x – 3) + 2

OpenStudy (anonymous):

for the graph of the line, we would need to see the possible graphs to answer

OpenStudy (anonymous):

this is the same as \[y = 4 x-4\]

OpenStudy (anonymous):

pick a line with slope 4 and y intercept \(-4\)

OpenStudy (anonymous):

c

OpenStudy (anonymous):

i think the first one works

OpenStudy (anonymous):

no C C has y intercept 5, that is if you posted them in order

OpenStudy (anonymous):

yeah did

OpenStudy (anonymous):

thank you so much i can i ever repay you

OpenStudy (anonymous):

then go with the first one, it is the only one that has y intercept \(-4\)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

thanks